# Novel Topological Insulators with Hybrid-order Boundary States

### v1

@@ Abstract
In the following we discuss the topology of this model in details. In AZ classification Chiu et al. (2016), ~$\tilde{\mathcal{H}}_{0}$ falls into class CII with $\mathcal{T}$, $\mathcal{C}$ and $\mathcal{S}$ hosting a $\mathbb{Z}_{2}$ topologyN (4). Since ~$\tilde{\mathcal{H}}_{0}$ is spin conservation with $\tilde{\Gamma}_{5}$-symmetry, we have $[\tilde{\Gamma}_{5},\tilde{\mathcal{H}}_{0}]=0$, where $\hat{\tilde{\Gamma}}_{5}=G_{300}$. We can define a $\mathbb{Z}_{2}$ number through the spin SCN N (5) to characterize the first-order topology with bulk-boundary correspondence even if $\mathcal{T}$ and $\mathcal{C}$ will be broken laterZhu et al. (2022). In addition, ~$\tilde{\mathcal{H}}_{0}$ also respects the following crystalline symmetries, including $\mathcal{I}$, and $\mathcal{M}_{i}$ leading to the combined symmetries $\mathcal{M}_{ij}$ and $\mathcal{R}_{i}\mathcal{T}=\mathcal{M}_{i}\mathcal{I}\mathcal{T}$ for $i=x,y,z,w$N (6).
@@
Figure 2: Schematic of the boundary states of $\mathcal{H}_{3}$ when $42<M<$ 4. (a-b)$b_{xw}=b_{yw}=b_{zw}>0$ and $c_{x}=c_{z}=0$, the system harbors both shifted first-order Dirac points located at the same positions as in Fig. 1(c) and second-order Fermi-arcs consisting of helical hinge states (colored in orange) of 3D SOTIs stacked along $k_{w}$ axis. (c-e) $b_{xw}=b_{yw}=b_{zw}>0$ and $c_{x}=c_{z}>0$. The system harbors both shifted first-order nodal lines and third-order Fermi-arcs consisting of zero-mode corner states of 3D TOTIs stacked along $k_{w}$ axis.
@@
We first consider the situation when $\Delta_{2}$ plays the role in, i.e., all $c_{i}=0$. $\Delta_{2}$ break $\mathcal{T}$, $\mathcal{C}$, but keep $\mathcal{S}$, $\mathcal{I}$ and thus the system is now falls into calss AIII but still hosts an unchanged spin SCN. The discussion of the second-order topology associated with the spin MFCNs N (8) is similar to the cases in $\mathcal{H}_{1}$ when only $b_{xw}$ and $b_{yw}$ surviveSM . To conveniently discuss third-order topology later, we focus on a TI phase with both first- and second-order boundary states with $b_{xw}=b_{yw}=b_{zw}$ for $M$ near $3$. As shown in Fig. 2(a-b), the system hosts the second-order helical Fermi-arcs where its 3D layers $\mathcal{H}_{2}^{k_{w}}$ with $k_{w}\in(-b_{xw},b_{xw})$ are SOTIs with helical states propagating along $\mathcal{M}_{x-y}$, $\mathcal{M}_{y-z}$, and $\mathcal{M}_{z-x}$ symmetric hingesN (1) when considering OBCs along $x$-, $y$- and $z$-directions. One can also define 3D mirror winding numbers to describe this second-order topology accordinglySM .
@@
Subsequently, we turn on $c_{x}=c_{z}$ being numeric small, $\mathcal{H}_{3}$ remains in class AIII harboring shifted 3D $\mathbb{Z}_{2}$ nodal lines Shao and Zhao (2018); Wang et al. (2020b); Ahn et al. (2018) on the 3D boundaries when assuming OBC along $x$ or $y$ directions [Fig. 2(c)]. However, helical hinge Fermi-arcs have been gapped due to $\Delta_{3}$ breaks $\tilde{\Gamma}_{5}$ ($\mathcal{M}_{x-y}$ and $\mathcal{M}_{y-z}$ are also broken), leading to third-order Fermi-arcs along $k_{w}$ axisSM , as shown in Fig. 2(c-e). In this case, even though the spin SCN is not well-defined, one can define a $\mathbb{Z}_{2}$ number as 4D generalization of Kane-Mele invariant Shao and Zhao (2018) since $\mathcal{H}_{3}$ remains respect $\mathcal{R}_{w}\mathcal{T}$-symmetryN (6). Alternately, we can define another topological number as $\nu=\nu_{0}-\nu_{\pi}~{}\text{mod}~{}2$ at $\mathcal{R}_{w}\mathcal{T}$-invariant lines $k_{w}=0,\pi$ with $\nu_{k_{w}}$ is a topological number of the 3D real TI with $\mathcal{S}$ Dai et al. (2021). Calculation results show that $\nu_{0}=1$ and $\nu_{\pi}=0$ for $M$ near $3$ SM . Meanwhile, $\mathcal{H}_{3}$ keeps $\mathcal{M}_{(z+x)yw}=\mathcal{M}_{z+x}\mathcal{M}_{y}\mathcal{M}_{w}$ (7), $\mathcal{M}_{z-x}$, and $\mathcal{I}$-symmetries. A non-trivial 1D mirror winding number $w_{1m}=2$ defined at $k_{x}=-k_{z}$ and $(k_{y},k_{w})=(0,0)$ associated with $\mathcal{M}_{(z+x)yw}$-symmetry SM guarantees the exact location of third-order Fermi-arc with $\mathcal{M}_{w}$-symmetry about $k_{w}=0$ consisting of a pair of zero-modes, accumulating at two $\mathcal{M}_{(z+x)y}$-symmetric [combined mirror reflection about $z=-x$ and $y=0$ planes] corners in a 3D real space cubic spanned in $(x,y,z)$. The bulk-corner correspondence is established here; $\mathcal{M}_{z-x}$ implies these two zero-modes live in $z=x$ plane; these Fermi-arc states are always $\mathcal{I}$-symmetric in $(x,y,z,k_{w})$ space; see Fig. 2(d-e). We emphasize that $\tilde{\Gamma}_{5}$ $(\mathcal{R}_{w}\mathcal{T})$ symmetry guarantees the existence of second (third)-order gapless Fermi arcs in class AIII, where these boundary modes are located at the positions determined by additional crystalline symmetries.
@@
Figure 3: Schematic of the boundary states of $\mathcal{H}_{3}$ when $42<M<$ 4 and $\delta>0$ being numeric small. (a) The system harbors both shifted first-order Weyl cones SM and the second-order chiral Fermi-arc hinge states. The green dashed squares denote the 3D TCIs being the unconventional phase transition points. (b) 3D non-trivial subsystems with fixing $k_{w}$ host second-order chiral hinge states.
@@
with a perturbation $\Delta_{4}=\delta(\cos k_{x}-\cos k_{y})\Gamma_{4}$ breaks $C_{4}^{zw}$ and $\mathcal{T}$ individually, but respects $C_{4}^{zw}\mathcal{T}$-symmetry CT . As shown in Fig. 3, $\Delta_{4}$ shifts the boundary Weyl cones along $k_{w}$-direction SM . Two Weyl cones located on different boundaries [e.g. $y(x)=-L/2$ and $y(x)=L/2$] are shifted along the same direction when OBC along $y$ ($x$)-direction is assumed, resulting in a second-order Fermi-arc at four hinges once we further open the boundary along $x$ $(y)$ axis. Except the unchanged $C_{2}$ for the first-order topolgy, we can define the so-called “rotational spin first Chern number” (RSFCN) at $C_{4}^{zw}\mathcal{T}$-invariant 2D momenta $(k_{x},k_{y})=(0,0),(\pi,\pi)$ which may be used to characterize the second-order topology SM . In addition, we can easily construct the spin conservation verion of $\mathcal{H}_{3}$ with eight-band as $\mathcal{H}_{4}(\bm{k})=\sigma_{3}\otimes\mathcal{H}_{3}$, which respects $C^{zw}_{4}\mathcal{T}$ and $\mathcal{S}$ and thus falls into class AIII. The system now hosts the first-order boundary Dirac modes and the second-order helical hinge Fermi-arcs. By setting $k_{z}=0$, we obtain the corresponding 3D unique TI. All these similar discussions about the second-order topology associated with additional crystalline symmetries are presented in Ref.SM .
@@
CTPTs and Novel TSMs.—By enlarging the parameters in $\Delta_{1}/\Delta_{2}$ of $\mathcal{H}_{1}$/$\mathcal{H}_{2}$, we can study the CTPT goes through a phase transition from this unique TI phase to a non-trivial nodal-line/sphere metallic phase and then finally becomes a trivial insulator. Such 4D nodal structures, including nodal-line, spin-nodal-line, and nodal-sphere, are characterized by the first Chern number (FCN) $C_{1}$, spin FCN $C_{1s}$, and 1D winding number $w_{1}$ respectively. The corresponding topological responses based on current works Wang et al. (2021); Ramamurthy and Hughes (2017); Hirsbrunner et al. (2023) are also discussed SM . On the other hand, we can obtain serveral new classes of TSMs with hybrid-order boundary Fermi-arcs by stacking the above TI phases along an extra spatial dimension $u$, as is commonly done, i.e., $M\rightarrow M^{\prime}=M-\cos k_{u}$, For instance, when $M=4$, $\mathcal{H}_{1}(M^{\prime})$/$\mathcal{H}_{1}^{k_{z}=0}(M^{\prime})$ hosts a pair of 5D/4D nodal surfaces/lines characterized by double charges $\nu_{D}=(C_{2}/w_{3},C_{1})$, where $w_{3}$ denotes the 3D winding number; $\mathcal{H}_{3}(M^{\prime})$/$\mathcal{H}_{3}^{k_{z}=0}(M^{\prime})$ hosts a pair of hybrid-order 5D/4D monopoles characterized by $C_{2}/w_{3}$. Two spin copies of these nodal-structures of $\mathcal{H}_{2}(M^{\prime})/\mathcal{H}_{4}(M^{\prime})$ can be also extended accordinglySM .
@@
Discussion and outlook.— In this work, we establish a valuable framework for understanding the origin of the distinctive properties of these novel TIs, demonstrating that these unique TIs can be seen as a higher-dimensional stacking effect of low-dimensional HOTIs undergoing UTPTs along an additional spatial dimension. In other words, each kind of $d$D HOTI corresponds to a $(d+1)$D unique TI phase for $d\geq 2$. The related nodal structures and their topological responses are also studied when $\mathcal{H}_{1}/\mathcal{H}_{2}$ becomes a gapless phase. Furthermore, the corresponding TSMs with novel nodal-structures can be obtained by stacking these TIs going through CTPTs along an extra spatial dimension. Note that these unique topological phases, especially for models with $d\geq 4$, could be experimental studied in sythetic matter, as could be realized in metamaterialsCheng et al. (2021); Ma et al. (2021); Chen et al. (2021a, b), in electric circuitsWang et al. (2020c, 2023), or in optical lattices with cold atomsPrice et al. (2015); Petrides et al. (2018); Lohse et al. (2018); Bouhiron et al. (2024), etc. In particular, we can utilize the cold-atom setup proposed in our previous work Zhu et al. (2022) to realize the 4D model ~$\tilde{\mathcal{H}}_{0}$ in Eq. (3), including the implementation of other models mentioned in this work in a similar manner SM . For the 3D model in $\mathcal{H}_{1}$ with $k_{z}=0$, $\mathcal{H}_{0}^{k_{z}=0}$ as a 3D strong TI in low-energy limit has been predicted Zhang et al. (2009) and realized Chen et al. (2009) in real materials, such as Bi2Se3 and Bi2Te3 families, one can consider to induce $\Delta_{1}$ term in these materials by magnetization or light-induced Floquet drivings N (10).
@@
We highlight that these novel topological phases can also be generalized to non-Hermitian Gong et al. (2018); Yao and Wang (2018) and Floquet systems Cayssol et al. (2013); Goldman and Dalibard (2014). The field-theoretical descriptions of these phases in the bulk is one of interesting directions. An intriguing and open question is the direct calculation of the effective action in the bulk. Take $\mathcal{H}_{1}$ as an example, the presence of shifted first-order boundary Weyl cones implies the possibility of an additional term beyond the $(4+1)$D Chern-Simons action in its 4D parent system. We provide a relevant discussion in the SM SM and will address it in the future. Overall, our work further open the door of investigating novel topological phases in the synthetic and real materials.
@@ Acknowledgements.
This work was supported by the Guangdong-Hong Kong Joint Laboratory of Quantum Matter, the NSFC/RGC JRS grant (Grant No. N  HKU774/21), the CRF (Grant No. C6009-20G) and GRF (Grants No. 17310622 and No. 17303023) of Hong Kong, Guangdong Provincial Quantum Science Strategic Initiative (GDZX2200001), and National Key R&D Program of China (Grant No. 2022YFA1405304).

### v3

@@ Abstract
In the following we discuss the topology of this model in detail. In AZ classification Chiu et al. (2016), ~$\tilde{\mathcal{H}}_{0}$ falls into class CII with $\mathcal{T}$, $\mathcal{C}$ and $\mathcal{S}$ hosting a $\mathbb{Z}_{2}$ topologyN (4). Since ~$\tilde{\mathcal{H}}_{0}$ is spin conservation with $\tilde{\Gamma}_{5}$-symmetry, we have $[\tilde{\Gamma}_{5},\tilde{\mathcal{H}}_{0}]=0$, where $\hat{\tilde{\Gamma}}_{5}=G_{300}$. We can define a $\mathbb{Z}_{2}$ number through the spin SCN N (5) to characterize the first-order topology with bulk-boundary correspondence even if $\mathcal{T}$ and $\mathcal{C}$ will be broken laterZhu et al. (2022). In addition, ~$\tilde{\mathcal{H}}_{0}$ also respects the following crystalline symmetries, including $\mathcal{I}$, and $\mathcal{M}_{i}$ leading to the combined symmetries $\mathcal{M}_{ij}$ and $\mathcal{R}_{i}\mathcal{T}=\mathcal{M}_{i}\mathcal{I}\mathcal{T}$ for $i=x,y,z,w$N (6).
@@
Figure 2: Schematic of the boundary states of $\mathcal{H}_{3}$ when $42<M<$ 4. (a-b)$b_{xw}=b_{yw}=b_{zw}>0$ and $c_{x}=c_{z}=0$, the system harbors both shifted first-order Dirac points located at the same positions as in Fig. 1(c) and second-order Fermi-arcs consisting of helical hinge states (colored in orange) of 3D SOTIs stacked along $k_{w}$ axis. (c-e) $b_{xw}=b_{yw}=b_{zw}>0$ and $c_{x}=c_{z}>0$. The system harbors both shifted first-order nodal lines and third-order Fermi-arcs consisting of zero-mode corner states of 3D TOTIs located at $(x,y,z)=(\mp L/2,\pm L/2,\mp L/2)$ stacked along $k_{w}$ axis.
@@
We first consider the situation when $\Delta_{2}$ plays the role in, i.e., all $c_{i}=0$. $\Delta_{2}$ breaks $\mathcal{T}$, $\mathcal{C}$, but keeps $\mathcal{S}$, $\mathcal{I}$ and thus the system is now falls into calss AIII but still hosts an unchanged spin SCN. The discussion of the second-order topology associated with the spin MFCNs N (8) is similar to the cases in $\mathcal{H}_{1}$ when only $b_{xw}$ and $b_{yw}$ surviveSM . To conveniently discuss third-order topology later, we focus on a TI phase with both first- and second-order boundary states with $b_{xw}=b_{yw}=b_{zw}$ for $M$ near $3$. As shown in Fig. 2(a-b), the system hosts the second-order helical Fermi-arcs where its 3D layers $\mathcal{H}_{2}^{k_{w}}$ with $k_{w}\in(-b_{xw},b_{xw})$ are SOTIs with helical states propagating along $\mathcal{M}_{x-y}$, $\mathcal{M}_{y-z}$, and $\mathcal{M}_{z-x}$ symmetric hingesN (1) when considering OBCs along $x$-, $y$- and $z$-directions. One can also define 3D mirror winding numbers to describe this second-order topology accordinglySM .
@@
Subsequently, we turn on $c_{x}=c_{z}$ being numeric small, $\mathcal{H}_{3}$ remains in class AIII harboring shifted 3D $\mathbb{Z}_{2}$ nodal lines Shao and Zhao (2018); Wang et al. (2020b); Ahn et al. (2018) on the 3D boundaries when assuming OBC along $x$ or $y$ directions [Fig. 2(c)]. However, helical hinge Fermi-arcs have been gapped due to $\Delta_{3}$ breaks $\tilde{\Gamma}_{5}$ ($\mathcal{M}_{x-y}$ and $\mathcal{M}_{y-z}$ are also broken), leading to third-order Fermi-arcs along $k_{w}$ axisSM , as shown in Fig. 2(c-e). In this case, even though the spin SCN is not well-defined, one can define a $\mathbb{Z}_{2}$ number as 4D generalization of Kane-Mele invariant Shao and Zhao (2018) since $\mathcal{H}_{3}$ remains respect $\mathcal{R}_{w}\mathcal{T}$-symmetryN (6). Alternately, we can define another topological number as $\nu=\nu_{0}-\nu_{\pi}~{}\text{mod}~{}2$ at $\mathcal{R}_{w}\mathcal{T}$-invariant lines $k_{w}=0,\pi$ with $\nu_{k_{w}}$ is a topological number of the 3D real TI with $\mathcal{S}$ Dai et al. (2021). Calculation results show that $\nu_{0}=1$ and $\nu_{\pi}=0$ for $M$ near $3$ SM . Meanwhile, $\mathcal{H}_{3}$ keeps $\mathcal{M}_{(z+x)yw}=\mathcal{M}_{z+x}\mathcal{M}_{y}\mathcal{M}_{w}$ (7), $\mathcal{M}_{z-x}$, and $\mathcal{I}$-symmetries. A non-trivial 1D mirror winding number $w_{1m}=2$ defined at $k_{x}=-k_{z}$ and $(k_{y},k_{w})=(0,0)$ associated with $\mathcal{M}_{(z+x)yw}$-symmetry SM guarantees the exact location of third-order Fermi-arc with $\mathcal{M}_{w}$-symmetry about $k_{w}=0$ consisting of a pair of zero-modes, accumulating at two $\mathcal{M}_{(z+x)y}$-symmetric [combined mirror reflection about $z=-x$ and $y=0$ planes] corners in a 3D real space cubic spanned in $(x,y,z)$. The bulk-corner correspondence is established here; $\mathcal{M}_{z-x}$ implies these two zero-modes live in $z=x$ plane; these Fermi-arc states are always $\mathcal{I}$-symmetric in $(x,y,z,k_{w})$ space [as seen in Fig. 2(d-e)]. We emphasize that $\tilde{\Gamma}_{5}$ $(\mathcal{R}_{w}\mathcal{T})$ symmetry guarantees the existence of second (third)-order gapless Fermi arcs in class AIII, where these boundary modes are located at the positions determined by additional crystalline symmetries.
@@
Figure 3: Schematic of the boundary states of $\mathcal{H}_{3}$ when $42<M<$ 4 and $\delta>0$ being numeric small. (a) The system harbors both shifted first-order Weyl cones SM and the second-order chiral Fermi-arc hinge states. The green dashed squares denote the 3D TCIs being the unconventional phase transition points. (b) 3D non-trivial SOTI hosts second-order chiral states located at four hinges.
@@
with a perturbation $\Delta_{4}=\delta(\cos k_{x}-\cos k_{y})\Gamma_{4}$ breaks $C_{4}^{zw}$ and $\mathcal{T}$ individually, but respects $C_{4}^{zw}\mathcal{T}$-symmetry CT . As shown in Fig. 3, $\Delta_{4}$ shifts the boundary Weyl cones along $k_{w}$-direction SM . Two Weyl cones located on different boundaries [e.g. $y(x)=-L/2$ and $y(x)=L/2$] are shifted along the same direction when OBC along $y$ ($x$)-direction is assumed, resulting in a second-order Fermi-arc at four hinges once we further open the boundary along $x$ $(y)$ axis. Except the unchanged $C_{2}$ for the first-order topology, we can define the so-called “rotational first Chern number” (RFCN) at $C_{4}^{zw}\mathcal{T}$-invariant 2D momenta $(k_{x},k_{y})=(0,0),(\pi,\pi)$ which may be used to characterize the second-order topology SM . In addition, we can easily construct the spin conservation verion of $\mathcal{H}_{3}$ with eight-band as $\mathcal{H}_{4}(\bm{k})=\sigma_{3}\otimes\mathcal{H}_{3}$, which respects $C^{zw}_{4}\mathcal{T}$ and $\mathcal{S}$ and thus falls into class AIII. The system now hosts the first-order boundary Dirac modes and the second-order helical hinge Fermi-arcs. By setting $k_{z}=0$, we obtain the corresponding 3D unique TI. All these similar discussions about the second-order topology associated with additional crystalline symmetries are presented in Ref.SM .
@@
CTPTs and Novel TSMs.—By enlarging the parameters in $\Delta_{1}/\Delta_{2}$ of $\mathcal{H}_{1}$/$\mathcal{H}_{2}$, we can study the CTPT goes through a phase transition from this unique TI phase to a non-trivial nodal-line/sphere metallic phase and then finally becomes a trivial insulator. Such 4D nodal structures, including nodal-line, spin-nodal-line, and nodal-sphere, are characterized by the first Chern number (FCN) $C_{1}$, spin FCN $C_{1s}$, and 1D winding number $w_{1}$ respectively. The corresponding topological responses based on current works Wang et al. (2021); Ramamurthy and Hughes (2017); Hirsbrunner et al. (2024) are also discussed SM . On the other hand, we can obtain serveral new classes of TSMs with hybrid-order boundary Fermi-arcs/surfaces by stacking the above TI phases along an extra spatial dimension $u$, as is commonly done, i.e., $M\rightarrow M^{\prime}=M-\cos k_{u}$, For instance, when $M=4$, $\mathcal{H}_{1}(M^{\prime})$/$\mathcal{H}_{1}^{k_{z}=0}(M^{\prime})$ hosts a pair of 5D/4D nodal-surfaces/lines characterized by double charges $\nu_{D}=(C_{2}/w_{3},\Delta C_{1})$, where $w_{3}$ denotes the 3D winding number; $\mathcal{H}_{3}(M^{\prime})$/$\mathcal{H}_{3}^{k_{z}=0}(M^{\prime})$ hosts a pair of hybrid-order 5D/4D monopoles characterized by $C_{2}/w_{3}$. Two spin copies of these nodal-structures of $\mathcal{H}_{2}(M^{\prime})/\mathcal{H}_{4}(M^{\prime})$ can be also extended accordinglySM .
@@
Discussion and outlook.— In this work, we have established a valuable framework in which these unique TIs can be viewed as a higher-dimensional stacking effect of low-dimensional HOTIs undergoing UTPTs along an additional spatial dimension, helping to elucidate their topological origin. These novel TIs, induced from first-order TIs, exhibit distinctive properties, including first-order topological boundary modes and higher-order Fermi arcs, which are presented for the first time in this work. In general, each kind of $d$D HOTI corresponds to a $(d+1)$D unique TI phase for $d\geq 2$. The related nodal structures and their topological responses are also studied when $\mathcal{H}_{1}/\mathcal{H}_{2}$ becomes a gapless phase. Furthermore, the corresponding TSMs with novel nodal-structures can be obtained by regularly stacking these TIs along an extra spatial dimension. Note that these unique topological phases, especially for models with $d\geq 4$, could be experimental studied in sythetic matter, such as metamaterials Cheng et al. (2021); Ma et al. (2021); Chen et al. (2021a, b), electric circuits Wang et al. (2020c); Yu et al. (2020); Wang et al. (2023), superconducting circuits Tan et al. (2018, 2021); Zhang et al. (2024), diamond with nitrogen vacancy center Kong et al. (2016); Yu et al. (2019); Chen et al. (2022), or cold atomic systems Zhu et al. (2006, 2007); Shao et al. (2008); Price et al. (2015); Goldman et al. (2014); Zhu et al. (2017); Petrides et al. (2018); Lohse et al. (2018); Bouhiron et al. (2024), etc. In particular, we can utilize the cold-atom setup proposed in our previous work Zhu et al. (2022) to realize the 4D model ~$\tilde{\mathcal{H}}_{0}$ in Eq. (3), including the implementation of other models mentioned in this work in a similar manner SM . For the 3D model in $\mathcal{H}_{1}$ with $k_{z}=0$, $\mathcal{H}_{0}^{k_{z}=0}$ as a 3D strong TI in low-energy limit has been predicted Zhang et al. (2009) and realized Chen et al. (2009) in real materials, such as Bi2Se3 and Bi2Te3 families, one can consider to induce $\Delta_{1}$ term in these materials by magnetization or light-induced Floquet drivings N (10).
@@
We highlight that these novel topological phases can also be generalized to non-Hermitian Shen et al. (2018); Gong et al. (2018); Yao and Wang (2018) and Floquet systems Cayssol et al. (2013); Goldman and Dalibard (2014); Nag et al. (2019); Salerno et al. (2020). The field-theoretical descriptions of these phases in the bulk is one of interesting directions in the future. Overall, our work further opens the door of investigating novel topological phases in the synthetic and real materials.
@@ Acknowledgements.
This work was supported by the National Key R&D Program of China (Grants No. 2024YFA1409300, No. 2022YFA1405304), the Guangdong-Hong Kong Joint Laboratory of Quantum Matter, the NSFC/RGC JRS grant (Grant No. N  HKU774/21), the CRF (Grant No. C6009-20G) and GRF (Grants No. 17310622 and No. 17303023) of Hong Kong, Guangdong Provincial Quantum Science Strategic Initiative (GDZX2200001), and Open Fund of Key Laboratory of Atomic and Subatomic Structure and Quantum Control (Ministry of Education).